OpenStudy (anonymous):

Given that f(x) = 2x+1 and g(x) = 3x-2, what value(s) of x are excluded from the domain of (f/g)(x)? I know how to set it up, but I don't understand the whole "exclusion" of the "domain"

Vocaloid (vocaloid):

 hint: the values excluded from the domain are: the x-values that will make the denominator equal to 0

OpenStudy (anonymous):

Yeah like when the = sign has a / across it, right? Would I divide 2x by 3x and 1 by 2?

Vocaloid (vocaloid):

nope, we can't do that because of these

Vocaloid (vocaloid):

don't change the function

OpenStudy (anonymous):

Oh right, so how would I do it?

Vocaloid (vocaloid):

set the denominator equal to 0 and solve for x

OpenStudy (anonymous):

so 3x-2=0?

Vocaloid (vocaloid):

right, so x = ?

OpenStudy (anonymous):

which would be 2/3

OpenStudy (anonymous):

Would I have to do the same thing to the numbers on the top? like 2x+1=0

Vocaloid (vocaloid):

nope, just the denominator

Vocaloid (vocaloid):

remember: when we divide two numbers, the numerator can be anything, but the denominator cannot be 0

OpenStudy (anonymous):

Ok, and then I would divide them together? or would I do KCF?

OpenStudy (anonymous):

Keep it change it flip it

Vocaloid (vocaloid):

no, you're done here, you found x = 2/3, that's your answer.

OpenStudy (anonymous):

Really?

OpenStudy (anonymous):

Oh wow.. Thank you!